If a lender charge a 9% nominal interest rate and expected inflation rate is a 4%, what is the difference between the real rate the lender received and the real rate the lender expected when the actual inflation ended up being 2%
Answer #1

Answer - Real interest rate = Nominal interest rate - Inflation

When inflation was 4 perc , real interest rate = 9-4 = 5%

Expected inflation rate when 2 perc , real interest rate = 9-2 = 7 perc.

Difference = 7- 5

= 2perc